OpenStudy (anonymous):

An auditorium earned $25,000 in sold−out concert ticket sales. Front section tickets cost $75 per seat and back section tickets cost $50 per seat. The number of front section seats is twice the number of back section seats. How many seats are in the front section?

OpenStudy (anonymous):

$75x+$50y=$25,000 and 2x=y

OpenStudy (anonymous):

is it right??hehe

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

actually its not

OpenStudy (anonymous):

@CandyCoatedRainDrops

OpenStudy (anonymous):

75(2B) + 50B = 25000 150B + 50B = 25000 200B = 25000 B = 125 F = 2B = 2(125) = 250

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

sorry x=2y....TT